### v1

@@ 1 Introduction
Among the forthcoming wide-field surveys, the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) [^1] https://www.lsst.org (Abell et al. 2009), the third generation of the South Pole Telescope (SPT-3G) [^2] https://astro.fnal.gov/science/cmbr/spt-3g/ (Benson et al. 2014), eROSITA [^3] https://www.mpe.mpg.de/eROSITA (Predehl et al. 2021), the Square Kilometre Array (SKA) [^4] https://www.skatelescope.org (Maartens et al. 2015), the Dark Energy Spectroscopic Instrument (DESI) [^5] https://www.desi.lbl.gov (Aghamousa et al. 2016), the Nancy Grace Roman Space Telescope [^6] https://roman.gsfc.nasa.gov (Spergel et al. 2015), and Euclid [^7] https://www.euclid-ec.org (Euclid Collaboration: Mellier et al. 2024), are poised to deliver measurements of summary statistics of the distribution of galaxy clusters with unprecedented statistical precision. These data sets will demand more accurate covariance models to exploit the cosmological information encoded in cluster clustering fully. The goal of this paper is, therefore, twofold. First, we investigate the cosmology dependence of the calibrated parameters that enter the covariance model for the cluster 2PCF. Second, we compare the parameters extracted from PINOCCHIO-based mocks with those calibrated on higher-fidelity $N$-body simulations. This analysis directly tests the robustness of the covariance modeling for future surveys such as Euclid, which is expected to deliver a photometric cluster catalog of unprecedented size and redshift coverage. Ultimately, a more reliable covariance prescription will improve cosmological constraints from cluster clustering and deepen our understanding of systematic uncertainties.
@@ 4.1 Resolution
Notice that the parameters $\alpha$ and $\beta$ in the covariance model have slightly different meanings and values compared to those fitted from the power spectrum. This is because the three covariance parameters also absorb the effects of missing higher-order terms in the covariance, partially losing their physical interpretation (Euclid Collaboration: Fumagalli et al. 2024). The agreement between these parameters will be discussed in Sect. 4.2.1.
@@ 4.3 Cosmological dependency
In Fig.8, we show the best-fit values for the covariance model from Eq.(4) as a function of cosmology. For a more realistic comparison, we perform the fit on lightcones, divided into different redshift bins. The parameter $\alpha$, while staying within the range $[-0.05,0.15]$—indicating small deviations from Poissonian shot noise—tends to increase with $\sigma_{8}$, suggesting a mild linear dependence on the amplitude of matter fluctuations. By contrast, $\beta$ (the bias rescaling) and $\gamma$ (the sub-leading shot-noise correction) do not show a strong trend with $\sigma_{8}$. However, both $\beta$ and $\gamma$ exhibit noticeable variations with redshift. Specifically, $\beta$ remains in the range [1.0, 1.5]1.01.5[1.0,\,1.5][ 1.0 , 1.5 ] for the different cosmologies but shifts slightly across redshift bins. At the same time, $\gamma$ is negative for most cosmologies, implying that the baseline Poisson noise may be an overestimate of the actual sampling noise in specific regimes but tends to decrease with increasing redshift. These results emphasize that even though the global amplitude of non-Poissonian effects and bias corrections is small, allowing for flexible parameters in the covariance model is essential to capture the redshift dependencies and subtle but non-trivial responses to variations in $\sigma_{8}$.
@@ 6 Conclusions
Overall, our results confirm that this semi-analytical covariance model can robustly capture the main sources of uncertainty in cluster clustering, providing the nuisance parameters are carefully tuned. In particular, approximate simulations such as PINOCCHIO can be employed effectively for such a calibration, enabling robust covariance estimates across diverse cosmological models and redshifts without incurring the computational expense of full $N$-body simulations.
@@
Authors’ Contributions
The authors, A. Fumagalli and T. Castro, share co-authorship of this paper. T. Castro conceptualized and produced the simulated data, while A. Fumagalli carried out the formal analysis of the simulated data. All authors contributed to interpreting the results and writing the manuscript. All authors have read and approved the final version of the paper.
@@ Acknowledgements.
It is a pleasure to thank Isabella Baccarelli, Fabio Pitari, and Caterina Caravita for their support with the CINECA environment. This work is supported by the Agenzia Spaziale Italiana (ASI) under - Euclid-FASE D Attivita’ scientifica per la missione - Accordo attuativo ASI-INAF n. 2018-23-HH.0, by the PRIN 2022 PNRR project ”Space-based cosmology with Euclid: the role of High-Performance Computing” (code no. P202259YAF), by the Italian Research Center on High-Performance Computing Big Data and Quantum Computing (ICSC), a project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, by the INFN INDARK PD51 grant, and by the PRIN 2022 project EMC2 - Euclid Mission Cluster Cosmology: unlock the full cosmological utility of the Euclid photometric cluster catalog (code no. J53D23001620006). AF acknowledges support by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311. AF acknowledges support from the Ludwig-Maximilians-Universität in Munich. The simulations were run partially on the Leonardo-Booster supercomputer as part of the Leonardo Early Access Program (LEAP) and under the ISCRA initiative. We acknowledge the CINECA award for the availability of high-performance computing resources and support. We acknowledge the use of the HOTCAT computing infrastructure of the Astronomical Observatory of Trieste – National Institute for Astrophysics (INAF, Italy) (see Bertocco et al. 2020; Taffoni et al. 2020).

### v2

@@ 1 Introduction
Among the forthcoming wide-field surveys, the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) [^1] https://www.lsst.org (LSST Science Collaborations: Abell et al. 2009), the third generation of the South Pole Telescope (SPT-3G) [^2] https://astro.fnal.gov/science/cmbr/spt-3g/ (Benson et al. 2014), eROSITA [^3] https://www.mpe.mpg.de/eROSITA (Predehl et al. 2021), the Square Kilometre Array (SKA) [^4] https://www.skatelescope.org (Maartens et al. 2015), the Dark Energy Spectroscopic Instrument (DESI) [^5] https://www.desi.lbl.gov (DESI Collaboration: Aghamousa et al. 2016), the Nancy Grace Roman Space Telescope [^6] https://roman.gsfc.nasa.gov (Spergel et al. 2015), and Euclid [^7] https://www.euclid-ec.org (Euclid Collaboration: Mellier et al. 2024), are poised to deliver measurements of summary statistics of the distribution of galaxy clusters with unprecedented statistical precision. These data sets will demand more accurate covariance models to exploit the cosmological information encoded in cluster clustering fully. The goal of this paper is, therefore, twofold. First, we investigate the cosmology dependence of the calibrated parameters that enter the covariance model for the cluster 2PCF. Second, we compare the parameters extracted from PINOCCHIO-based mocks with those calibrated on higher-fidelity $N$-body simulations. This analysis directly tests the robustness of the covariance modeling for future surveys such as Euclid, which is expected to deliver a photometric cluster catalog of unprecedented size and redshift coverage. Ultimately, a more reliable covariance prescription will improve cosmological constraints from cluster clustering and deepen our understanding of systematic uncertainties.
@@ 4.1 Resolution
Notice that the parameters $\alpha$ and $\beta$ in the covariance model have slightly different meanings and values compared to those fitted from the power spectrum. This is because the three covariance parameters also absorb the effects of missing higher-order terms in the covariance, partially losing their physical interpretation (Euclid Collaboration: Fumagalli et al. 2024). The agreement between these parameters will be discussed in Sect. 4.2.
@@ 4.3 Cosmological dependency
In Fig.8, we show the best-fit values for the covariance model from Eq.(4) as a function of cosmology. For a more realistic comparison, we perform the fit on lightcones divided into different redshift bins. The parameter $\alpha$, while staying within the range $[-0.05,0.15]$—indicating small deviations from Poissonian shot noise—tends to increase with $\sigma_{8}$, suggesting a mild linear dependence on the amplitude of matter fluctuations. By contrast, $\beta$ (the bias rescaling) and $\gamma$ (the sub-leading shot-noise correction) do not show a strong trend with $\sigma_{8}$. However, both $\beta$ and $\gamma$ exhibit noticeable variations with redshift. Specifically, $\beta$ remains in the range [1.0, 1.5]1.01.5[1.0,\,1.5][ 1.0 , 1.5 ] for the different cosmologies but shifts slightly across redshift bins. At the same time, $\gamma$ is negative for most cosmologies, implying that the baseline Poisson noise may be an overestimate of the actual sampling noise in specific regimes but tends to decrease with increasing redshift. These results emphasize that even though the global amplitude of non-Poissonian effects and bias corrections is small, allowing for flexible parameters in the covariance model is essential to capture the redshift dependencies and subtle but non-trivial responses to variations in $\sigma_{8}$.
@@ 6 Conclusions
Overall, our results confirm that this semi-analytical covariance model can robustly capture the main sources of uncertainty in cluster clustering, provided the nuisance parameters are carefully tuned. In particular, approximate simulations such as PINOCCHIO can be employed effectively for such a calibration, enabling robust covariance estimates across diverse cosmological models and redshifts without incurring the computational expense of full $N$-body simulations. Furthermore, the demonstrated success of PINOCCHIO in reproducing the cluster covariance motivates the extension of our analysis to a multi-probe context in future studies. Importantly, assessing its capability to estimate covariances for additional cluster-based summary statistics, such as cluster shear correlation and cluster galaxy correlation, will allow a more comprehensive evaluation of approximate methods in various cosmological applications.
@@ Acknowledgements.
The authors A. Fumagalli and T. Castro share co-authorship of this paper. T. Castro conceptualized and produced the simulated data, while A. Fumagalli carried out the formal analysis of the simulated data. All authors contributed to interpreting the results and writing the manuscript. All authors have read and approved the final version of the paper. It is a pleasure to thank Isabella Baccarelli, Fabio Pitari, and Caterina Caravita for their support with the CINECA environment. This work is supported by the Agenzia Spaziale Italiana (ASI) under - Euclid-FASE D Attivita’ scientifica per la missione - Accordo attuativo ASI-INAF n. 2018-23-HH.0, by the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 1409 published on 14.9.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title ”Space-based cosmology with Euclid: the role of High-Performance Computing” – CUP J53D23019100001 - Grant Assignment Decree No. 962 adopted on 30/06/2023 by the Italian Ministry of Ministry of University and Research (MUR); by the Italian Research Center on High-Performance Computing Big Data and Quantum Computing (ICSC), a project funded by the European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, by the INFN INDARK PD51 grant, and by the PRIN 2022 project EMC2 - Euclid Mission Cluster Cosmology: unlock the full cosmological utility of the Euclid photometric cluster catalog (code no. J53D23001620006). AF acknowledges support by the Excellence Cluster ORIGINS, which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311. AF acknowledges support from the Ludwig-Maximilians-Universität in Munich. The simulations were run partially on the Leonardo-Booster supercomputer as part of the Leonardo Early Access Program (LEAP) and under the ISCRA initiative. We acknowledge the CINECA award for the availability of high-performance computing resources and support. We acknowledge the use of the HOTCAT computing infrastructure of the Astronomical Observatory of Trieste – National Institute for Astrophysics (INAF, Italy) (see Bertocco et al. 2020; Taffoni et al. 2020).